A hoop, a solid cylinder, a solid sphere, and a thin spherical shell each has the same mass of 2.68 kg and the same radius of 0.204 m. Each is also rotating about its central axis with an angular speed of 37.0 rad/s. What is the magnitude of the angular momentum of each object? (Enter your answers in kg · m2/s.)

(a) hoop kg · m2/s

(b) solid cylinder kg · m2/s

(c) solid sphere kg · m2/s

(d) thin, spherical shell kg · m2/s
Answer #1

(1) L- IL - MR & =(2.68)(0.204) * 37 = 0.11153437 = 4.1260 kg ml L - IX (2) - 1 MR²2 = 1x4.1266 22.063 169 m/s IL (3) L - 2 M

Please rate the work if you are satisfied with the solution or else kindly ask your doubts or queries in the comment section below